Assume the following data for Cable Corporation and Multi-Media Inc.
           Cable
           Corporation    Multi-Media Inc.
Net income    $    35,800        $ 128,000   
Sales            346,000       2,700,000   
Total assets        449,000       929,000   
Total debt        201,000    523,000   
Stockholders' equity    248,000    406,000   

a-1. Compute return on stockholders’ equity for both firms

a-2. Which firm has the higher return

b. Compute the following additional ratios for both firms

net income sales (%)
net income/total assets (%)
sales/total assets (times)
debt/total assets (%)
Answer #1

a-1]

ROE = net income / stockholders equity

Cable Corp = 35,800 / 248,000 = 14.44%

MultiMedia = 128,000 / 406,000 = 31.53%

a-2]

MultiMedia has the higher return

b]

net income / sales

Cable Corp = 35,800 / 346,000 = 10.35%

MultiMedia = 128,000 / 2,700,000 = 4.74%

net income /total assets

Cable Corp = 35,800 / 449,000 = 7.97%

MultiMedia = 128,000 / 929,000 = 13.78%

sales /total assets

Cable Corp = 346,000 / 449,000 = 0.77 times

MultiMedia = 2,700,000 / 929,000 = 2.91 times

debt /total assets

Cable Corp = 201,000 / 449,000 = 44.77%

MultiMedia = 523,000 / 929,000 = 56.30%